T-474 phone conference

Thursday 17 august 2006, 09:00 SLAC time, 17:00 UK time

Phone number : +1 510-665-5437
Meeting code : 9115

hardware points


analysis points

  • Raw waveforms from TDS oscilloscope, tds2root, see T474 eLog entry 49
  • State of the ESA data
    • filling gaps of july data taking, first processing run
    • reprocessing with latest version to include VSAM data, timestamp + 'upgrade' to the filtering algoritm to account for small residual phase dependence
    • set of routines for easy dataanalysis... libRooEsa needs work !!
    • stability data analysis
    • stability check + analysis of BPMCALI macro data
    • orbit fitting
  • Reprocessing of T-480 runs, list with required runs from july run
  • Checks + status of libespec library (Mark), see CVS server
    Restructuring of the library to make it easy to work with and stable, see elog entry
    Still large todo list, which some items that have to be checked/done asap.
  • Documentation page for contents of the ROOT ntuples


Minutes of the 17 August 2006 T474 Phone conference

- As Alexey, Stewart and Ray were not able to phone, most of the
  in depth hardware discussions on the calibration system, the new BPM
  electronics and Ray's proposal for the magnets, will be postponed to
  next week.

- Mike reported on the SLAC budget review. Bottom line is that he is
  optimistic about getting funding through 2007, however there is some
  uncertainty about 2008. There is appearantly a big push to do things
  in 2007. This would mean extra pressure on getting the BPM prototype
  installed in the end station for the first run next year.
  As the UCL workshop, which is preparing an aluminium test BPM, is
  delayed, this looks uncertain. We will know more from the workshop
  early next week. Alexey will be able to report on that in more
  detail next thrusday.

- As to the magnets, a couple of items have to be purchased
   o probably need a chiller for the cooling water for the magnets
   o need to find a power supply for the magnets, depending on how we
     want to feedback : on NMR probes or on current ? Mike thinks that
     the company that is making the NMR probes also has some power
     supplies that use hardware feedback.
  Everyone agrees that it makes more sense to proceed with the 10D37
  magnets. Mike had some discussions on the magnets as to how precise
  we want to know the integrated B.dl. Do we need to know it to 10^-5,
  as there is probably nothing to compare it with, so this might be
  difficult. It is probably more interesting to probe for stability !

- Bino briefly ran over the plots that were in Alexey's eLog, a
  summary with explanations from Alexey is available (see agenda).

- Yury will the geophone measurements in the end statio ntogether with Mahsa.

- Bino quickly reported on the reprocessing of the data of the raw
  scope. These data are basically understood by Zenghai and Chris,
  however the equidistant little peaks in the FFT remain not
  understood fully, although it looks like an effect of the scope

- Bino will reprocess/calibrate the missing runs that were taking
  during the online datataking, including the T-480 runs as requested
  by Johnny. Yury is updating the code to include the VSAM data and
  have a second filter suppressing the residual 2omega frequency
  peak. We will need a new round of calibration/optimisation

- Bino reported briefly on the status of libespec, more info is in the

- We decided to have a phone conference next week, Yury has already
  booked a slot, info will be sent to the list.

- Yury stated that in future, with the new term starting, phone
  conferences on thursday might be a bit inconvenient, there is a
  proposal to shift them to friday, will send this around and discuss
  by mail.
back to main